# Optimal Execution in Intraday Energy Markets under Hawkes Processes with Transient Impact

### v1

@@ 4.2 Benchmark strategies
and at maturity $X_{T}^{TWAP}=0$. So that the agent’s trading rate is constant (constant velocity) and $\xi_{t}^{TWAP}\equiv\frac{X_{0}}{T}$ for all $t\in[0,T]$. It has been shown in the seminal work of Almgren and Chriss [5] that under the martingale assumption on the execution price dynamics, linear permanent impact, and excluding risk-aversion such a strategy is optimal in terms of minimizing the expected quadratic execution costs.
Since we are considering the framework where apart from the permanent and the temporary market impact component, we also have a transient component in the impact of the trading rate, we will also consider as a strategy the one derived by Obizhaeva and Wang in [31].
@@
One of the most commonly used benchmark strategies in algorithmic trading is the Volume-Weighted Average Price (VWAP). It is straightforward to compute ex post, making it a popular performance metric: if execution prices remain close to the VWAP, then the trader has effectively matched the volume-weighted average price over the execution window. However, implementing a VWAP strategy in real time is challenging because the intraday distribution of total traded volume is unknown in advance. Although volume in intraday markets often increases toward gate closure [26], substantial variability persists both intraday and across different days.
@@
Targeting a VWAP execution schedule is difficult due to the random nature of traded volume throughout the day. One practical approach is to submit orders at a rate proportional to the ongoing market order flow, approximating a fraction of the total trading activity [12]. Because we model trading intensity, a VWAP-like strategy can be formulated by partitioning the trading horizon into minute buckets and distributing trades in proportion to the historical (or forecast) fraction of volume within each bucket.
@@
Figure 10 shows an example of the resulting trading trajectory compared with TWAP and the Hawkes-derived optimal strategy, assuming $X_{0}=250\,\mathrm{MWh}$ for liquidation across four representative hourly contracts (15h–18h). The paths illustrate how the VWAP strategy seeks to align trading with the estimated intraday volume distribution, while TWAP divides the trading horizon evenly, and the Hawkes-based method exploits predicted intensity dynamics for potentially lower costs.
@@
Figure 10: Trading trajectories (Residual position) for the optimal execution strategy (simulations), TWAP, and VWAP for the last 8 trading hours assuming a total of $X_{0}=250$ MWh to be liquidated. The simulations are based on parameters fitted for the 15h, 16h, 17h, and 18h products.
One can easily understand that different trading algorithms result in different trading trajectories. Having presented the mathematical formulation of our optimal execution strategy as well as the benchmark strategies that we consider for the backtesting of the optimal strategy, we showcase a graph with the trading position of each strategy to sharpen the understanding of the reader using some numerical results. The easiest and most straightforward way to understand volume allocation behaviour and the metaorder execution problem (in terms of a liquidation problem) is by inspecting a trade path plot. Figure (10) showcases a MC simulation of the optimal trading trajectory, the VWAP and TWAP strategy. Considering that positions 200-300 MWh apply to many market participants and depict an industry-wide optimal execution problem, we consider that the trader wants to liquidate a total of 250 MWh in the ID market in a time window of T = 8h. We simulate the optimal strategy using a total of $N_{sims}=1000$.
For this simulation we consider the calibrated parameters for the 15h, 16h, 17h, and 18h products as well as the empirical volumes calculated for this hourly products. As we can see from the figure the VWAP strategy starts trading with slower trading rates and leaves a higher position open in the early/mid-trading phases. Just before the switch from pan-European XBID to local trading, we can see that positions are closed more aggressively leaving a smaller position in the last trading hour. Finally, the optimal execution strategy follows the trading trajectory of the VWAP strategy, however being even more passive at the start of the trading phase and being aggressive during the end. For the last trading hour, we can see how the strategy resembles even more than one of a VWAP execution type. Taking a look at the trading trajectories of the optimal strategy $X_{t}^{*}$ and the $X_{t}^{VWAP}$ these trading paths avoid both the higher BAS in the early phase of the trading (far from delivery) but also the high BAS levels in local trading. As we have already described Figure 9 showcases the mean instantaneous market impact coefficient with respect to the time to gate closure, showcasing the potential profitability of the optimal execution strategy, when our objective is to minimize the transaction costs of the liquidation strategy.
@@ 4.3 Backtesting and Transaction Cost Analysis
$\mathcal{D}^{\text{Strat}}$ $=\sum_{t=1}^{T}\xi_{t}^{\text{Strat}}\left(\sum_{1\leq t^{\prime}<t}G(t-t^{\prime})\xi_{t^{\prime}}^{\text{Strat}}\right)$ (37)
$+\sum_{t=1}^{T}\frac{\xi_{t}^{\text{Strat}}s_{t}}{2}$ (38)
where we have used that $p_{t}=m_{t}+s_{t}/2$, with $s_{t}/2$ the half-spread. We have to note that in our modeling framework the sign of the metaorder is uncorrelated with the order flow from the rest of the market. For optimal trade execution under endogenous order flow we refer the reader to the following work [15]. Equation (37) shows that the expected execution shortfall is partly due to the spread and partly due to the impacts of past trades. Our goal is to use the expected shortfall as our cost function to back test the optimal execution strategy to realized market data, by minimizing this cost functional. We will also consider the simplified case where the transient market impact component boils down to the instantaneous impact only which we have estimated using the procedure described in Section 3.2. For the half-spread cost we use the empirical time-weighted half-spread directly obtained from the LOB data as described in 3.2.
We will implement the optimal execution strategy for each hourly product using the estimated parameters we have derived, both for the Hawkes process which models the arrival rates of trades (buy/sell). We are going to distinguish between the univariate and bivariate calibration using the obtained parameters we presented in Tables 1 for the univariate case and 5 for the bivariate case accordingly. The transaction cost analysis results, will be devoted to each hourly product separately. Our goal is to test whether the optimal strategy allows for cost reduction when we are considering the liquidation of a fixed position of total $X_{0}=200-300$ MWh. The relative performance measures with respect to the optimal strategy is given by $r_{\text{Bench}}=(\mathcal{D}^{\text{Bench}}-\mathcal{D}^{\text{Strat}})/\mathcal{D}^{\text{Bench}}$. Starting from the univariate case, where the arrival of buy and sell market orders are independent, Figure (11) showcases the relative cost improvements (in percentage points) for each hourly trading product in comparison to the TWAP and VWAP strategies. The relative improvements in terms of cumulative execution costs are on average positive, indicating a clear cost reduction when the trader chooses to liquidate their position using the optimal execution strategy.
@@
Examining the average relative cost improvements associated with the optimal strategy and the two benchmarks for each maturity, we observe that the optimal strategy achieves cost reduction for each product. Specifically, the cost reduction is very significant for the first hourly trading products (1h-4h) and then stabilizes for the rest of the trading hours, exhibiting a slight decrease for the mid-day trading products (12h-16h).
Considering the seasonality of liquidity across each trading product, as shown in Figures (2), there is a clear pattern in the associated liquidity of each hourly product. The most liquid products are associated with mid-day (noon) hours, while the first hourly trading products are the least liquid ones. This suggests that the amount of cost savings is negatively correlated with the average traded volume observed for each product. Therefore, comparatively less volatile, more liquid trading products at lower tick sizes may benefit less from the optimal trading strategy compared to the less liquid products, where the improvement is more pronounced.
Looking closely at the path-dependent costs (rather than cumulative ones) associated with each strategy, Figures 33 showcase the relative cost improvement of the optimal strategy compared to the two benchmarks TWAP/VWAP, respectively. A positive value in the surface plot indicates higher costs for the TWAP/VWAP strategy, and vice versa. The results are plotted in a surface plot to illustrate how the relative improvement behaves for each trading hour.
In both cases, the optimal strategy incurs smaller costs at the beginning of the trading activity, which is expected since the speed of order execution is slower at the start of the trading session for the optimal strategy, accelerating as time to maturity shrinks. One observation from Figures 33 is that right after the market decoupling (local-trading regime), the mean-relative cost improvement of the optimal strategy seems to decrease. This indicates that the strategy does not manage to close a significant portion of the position before the market decoupling, resulting in higher execution costs due to the high BAS levels in the local trading regime. However, the strategy seems to recover, and the final relative execution costs with respect to the TWAP and VWAP strategies at maturity $T=8$ are positive.
@@
Figure 12: Visual comparison of relative cost improvements across maturities.
@@
Looking at the mean relative cost improvements of the bivariate model compared to the VWAP, TWAP and the univariate model the following table showcases that the results compared to the benchmark strategies are again significant. However, compared to the univariate case, where we have only the self-excitation parameters in the exponential kernel, the results vary across each hourly trading product, indicating that there is no clear improvement when considering the full kernel for the modeling of the order flow of buy and sell MOs.
@@
Figure 13: Visual comparison of relative cost improvements across maturities under self- and cross-excitation.
@@ 5 Conclusion
There are several potential extensions to the methodology and modeling framework presented in this work. One possible extension is to model multiple contracts simultaneously rather than focusing on a single contract. Our current one dimensional optimal execution model does not capture the dependence between different maturities, which is crucial for valuing assets with payoffs depending on several maturities. This could be addressed by employing a multivariate transient price impact kernel which models how trades for a given trading hourly product would (directly or indirectly) impact the price of another product.
One could also formulate an optimal execution problem under a cross-impact propagator model, where in intra-day markets overlapping sessions exist for different delivery days. Given the strong interdependencies among products, incorporating cross-product and cross-impact effects may offer deeper insights into price formation in intraday electricity markets.
@@ Acknowledgement
Konstantinos Chatziandreou gratefully acknowledges the support through Statkraft Trading GmbH. Konstantinos Chatziandreou is grateful to Daniel Gruhlke and Simon Hirsch for many helpful discussions.This work contains the author’s opinion and does not necessarily reflect Statkraft’s position. The authors declare no conflict of interest.
@@ Appendix A MLE for Hawkes Parameters
[…]
@@
[…]
@@ Appendix B Goodness of fit analysis
In this section we aim to describe how we selected the best model from the estimated point process models describing the order-flow. We are first going to introduce a general theory for the goodness of fit analysis of point processes. To this point let the compensator (or the cumulative intensity process) of a point process $N_{t}$ be defined by the following integral of the conditional intensity
[…]
By using the compensator, the technique of random time change transforms a point process with intensity $\lambda(t)$ into a standard Poisson process. The underlying rationale is that if the proposed point process model aligns well with the observed data, the time-changed process should exhibit the defining characteristics of a standard Poisson process. The following theorem, often referred to as the "random time change theorem," underpins this principle and serves as the foundation for the goodness-of-fit analysis we will present.
Theorem B.1.
(Papangelou’s Random Change time Theorem see [16] Theorem 7.4.I) Let $\left(\Omega,\mathcal{F},\mathbb{P},\mathbb{F}\right)$ be a filtered probability space and let $N_{t}$ be a simple point process, which is adapted to the filtration $\mathbb{F}$ and has a bounded and strictly positive intensity $\lambda(t)$. Moreover suppose that the compensator $\Lambda(t)$ with $\Lambda(t)$ defined as above is $\mathbb{P}$ a.s not bounded. Define the random time change $t\mapsto\tilde{t}=\Lambda(t)$. Then the process
$\tilde{N}(\tilde{t})=N(\Lambda^{-1}(\tilde{t}))$ (43)
is a standard Poisson process.
The core idea of this proof is to demonstrate that, after applying the appropriate random time change, the interarrival times become independent and identically distributed (i.i.d.) exponential random variables with a mean of one. Consider a point process $N_{t}$ whose events occur at times $\tau_{n}$. According to the theorem, the transformed arrival times $\tilde{\tau}_{n}=\Lambda(\tau_{n})$ correspond to the arrival times of a unit-rate Poisson process. As a result, the transformed interarrival times
[…]
@@
B.1 Goodness of Fit results - Comparison Piecewise and Spline approximation
@@
Starting from the univariate case, which is equivalent to the model where we have a zero component in the non-diagonals of the exponential kernel matrix of the Hawkes process we examine the two different choices of the baseline function. This model essentially boils down to the model where the order-flow is uncorrelated, meaning that buy market orders only excite orders of the same sign. The first and the most naive one would be to look into hourly blocks of the LOB and for each of these blocks to fit the underlying Hawkes process in order to account for the regime switch of the baseline intensity. Of course, this naive approach is also time-consuming. Considering a total of one year of back-test data, along with the calibration of each of the 24 products and the last 8 tradable hours, this approach does not seem appealing at first sight. Hence the second approach would be to consider an exponential baseline function. However, as we described above we are now going to consider a new approach by non-parametrically estimating the baseline intensity using the idea originally introduced in the paper [13] to account for the empirical finding motivated by Figure 2, which describes the drop of trading activity during the local trading hours and the sharp increase towards delivery for the first trading hours. We can see from the following Figure 14 that both the piecewise approximation  [^1] Looking into hourly blocks of the LOB and fitting the Hawkes parameters by assuming a constant baseline intensity for each hourly block as well as the non-parametric approximation of the baseline intensity seem to capture the increase of the market activity for both the sell and buy side of the LOB. However, it is clear from the graph that the median baseline intensities estimated using the non-parametric approach seem to better fit the drop in the market activity immediately after the start of the local trading regime, for both sides of the LOB; which is also in agreement with the empirical intensities presented in Figures 2. In Figure 14 we showcase the result only for the hourly products (16h - 18h). The total number of basis points used for the estimation of the baseline intensity is $N_{basis}=10$. One more nice feature that we observe is that on average the shape of the baseline function for each delivery hour seems to be close. This is also showcased in Figure 5 where we present the calibrated baseline intensity for each side of the LOB and each hourly product.
@@
In order to evaluate the performance of each model we perform a Goodness-of Fit analysis. For the comparison of the models, we use the calibrated models for each maturity (24 in total) between the dates (2023-01-26 until 2023-12-03 a total of 311 days). Hence we have a total of 7464 models for each different approximation of the baseline intensity. Since both models have been calibrated using a Maximum Likelihood method, one way to compare the fitting performance of each model is by using a Likelihood ratio test. However, in this work, we simply use a more informal inspection of the differences in the maximum log-likelihood values of the model with a baseline piecewise constant function and the model with a spline approximation of the baseline function. Since we are calibrating each model independently we showcase the differences for each hourly contract and each daily calibration. The following Figure showcases the box plots of the difference for each hourly product and each orderbook side (Buy/Sell). For both sides and each hourly the Maximized value of the log-likelihood function for the model with piece wise constant baseline function is on average lower than the spline approximation model, which is attributed to the positive differences of the maximum likelihood values showcased in the Box plots 16 - 17  [^2] We note that the boxplots include all fitted models; (2023-01-26 until 2023-12-03 a total of 311 days) for buy and sell market orders respectively..
@@
In a similar fashion we compare the AIC differences for each hourly product. Recall that in comparison to the log-likelihood tests the AIC scores indicate the best-performing model when the score is smaller. From the following figure, we can again see that the AIC score of the spline model is smaller (since we are testing the differences between (AIC Piecewise- AIC Spline) and we can see a positive difference between the two) for each calibrated trading hour. This indicates that the performance of the spline fit on the baseline intensity of the Hawkes process provides a better fit to the data. This result is obtained for both sides of the LOB (Buy/Sell) (See Figures 18-19).
@@
[…]
This enables us to validate the estimated model using both the Anderson-Darling Test as well as the Wasserstein distance on an exponential distribution with parameter $\lambda=1$ (see also [10]). The following figures depict the Wasserstein distances for buy and sell market orders respectively and include once more all the fitted models (2023-01-26 until 2023-12-03). We only showcase the results for three maturities due to a lack of space, however, the rest of the calibrated hourly products can be found in Appendix C. Looking at the Buy/Sell side, the median Wasserstein distance for the model with the spline approximation of the baseline intensity is significantly lower than for the piece wise approximation (see Figures 20-21). Additionally, looking at the inter quantile range (IQR) of the distance for the case of the spline model it is considerably lowe than the piecewise approximation. The Anderson-Darling test showcases that the fit to a unit exponential is better on both sides of the LOB since the values of the test statistic are lower in both cases. Finally, we examine the independence for the time-changed interarrival times using the Ljung-Box test. The results of the p-values plot in Figures (24-25), again reveal that since the mean of the p-values of all fitted models is not below $0.05$, which is the selected significance level, both tests for the spline and piecewise models agree that you can not reject the null of no auto-correlation between the series and each of its first 100 lags with > 95% confidence level. However testing the number of times that the null hypothesis was rejected between all fitted models (days), and examining the difference between the spline and the piece wise model, we can conclude that this number is smaller for the spline model, indicating that most of the times the time-changes inter arrival times of the spline model are indeed serially uncorrelated.
@@
Figure 26: Calibrated values $\alpha$ parameters for Hawkes process exponential excitation kernel (Buy/Sell side) and (Spline/Piece wise baseline approximation)
Figure 27: Calibrated values $\beta$ parameters for Hawkes process exponential excitation kernel (Buy/Sell side) and (Spline/Piece wise baseline approximation)

### v2

@@ 4.2 Benchmark strategies
and at maturity $X_{T}^{TWAP}=0$. So that the agent’s trading rate is constant (constant velocity) and $\xi_{t}^{TWAP}\equiv\frac{X_{0}}{T}$ for all $t\in[0,T]$. It has been shown in the seminal work of Almgren and Chriss [almgren2000optimal] that under the martingale assumption on the execution price dynamics, linear permanent impact, and excluding risk-aversion such a strategy is optimal in terms of minimizing the expected quadratic execution costs.
Since we are considering the framework where apart from the permanent and the temporary market impact component, we also have a transient component in the impact of the trading rate, we will also consider as a strategy the one derived by Obizhaeva and Wang in [Obizhaeva2005].
@@
One of the most commonly used benchmark strategies in algorithmic trading is the Volume-Weighted Average Price (VWAP). It is straightforward to compute ex post, making it a popular performance metric: if execution prices remain close to the VWAP, then the trader has effectively matched the volume-weighted average price over the execution window. However, implementing a VWAP strategy in real time is challenging because the intraday distribution of total traded volume is unknown in advance. Although volume in intraday markets often increases toward gate closure [KZ20], substantial variability persists both intraday and across different days.
@@
Targeting a VWAP execution schedule is difficult due to the random nature of traded volume throughout the day. One practical approach is to submit orders at a rate proportional to the ongoing market order flow, approximating a fraction of the total trading activity [CJ16]. Because we model trading intensity, a VWAP-like strategy can be formulated by partitioning the trading horizon into minute buckets and distributing trades in proportion to the historical (or forecast) fraction of volume within each bucket.
@@
Figure 8 shows an example of the resulting trading trajectory compared with TWAP and the Hawkes-derived optimal strategy, assuming $X_{0}=250\,\mathrm{MWh}$ for liquidation across four representative hourly contracts (15h-21h). The paths illustrate how the VWAP strategy seeks to align trading with the estimated intraday volume distribution, while TWAP divides the trading horizon evenly, and the Hawkes-based method exploits predicted intensity dynamics for potentially lower costs.
@@
(e) residual position for the 19h product.
(f) Residual position for the 21h product.
Figure 8: Trading trajectories (Residual position) for the optimal execution strategy (simulations), TWAP, and VWAP for the last 8 trading hours assuming a total of $X_{0}=250$ MWh to be liquidated. The simulations are based on parameters fitted for the 15h, 16h, 17h, 18h, 19h, 21h products.
One can easily understand that different trading algorithms result in different trading trajectories. Having presented the mathematical formulation of our optimal execution strategy as well as the benchmark strategies that we consider for the backtesting of the optimal strategy, we will present some numerical results. The most straightforward way to understand volume allocation behaviour and the metaorder execution problem (in terms of a liquidation problem) is by inspecting a trade path plot. Figure 8 showcases a MC simulation of the optimal trading trajectory, the VWAP and TWAP strategy. Considering that positions 200-300 MWh apply to many market participants and depict an industry-wide optimal execution problem, we consider that the trader wants to liquidate a total of 250 MWh in the ID market in a time window of T = 8h. We simulate the optimal strategy using a total of $N_{sims}=1000$.
For this simulation we consider the calibrated parameters for the 15h, 16h, 17h, 18h, 19h and 21h products as well as the empirical volumes calculated for this hourly products. As we can see from the figure the VWAP strategy starts trading with slower trading rates and leaves a higher position open in the early/mid-trading phases. Just before the switch from pan-European XBID to local trading, we can see that positions are closed more aggressively leaving a smaller position in the last trading hour. Finally, the optimal execution strategy follows the trading trajectory of the VWAP strategy, however being even more passive at the start of the trading phase and being aggressive during the end. For the last trading hour, we can see how the strategy resembles even more than one of a VWAP execution type. Taking a look at the trading trajectories of the optimal strategy $X_{t}^{*}$ and the $X_{t}^{VWAP}$ these trading paths avoid both the higher BAS in the early phase of the trading (far from delivery) but also the high BAS levels in local trading. As we have already described Figure 7 showcases the mean instantaneous market impact coefficient with respect to the time to gate closure, showcasing the potential profitability of the optimal execution strategy, when our objective is to minimize the transaction costs of the liquidation strategy.
@@ 4.3 Backtesting and Transaction Cost Analysis
$\mathcal{D}^{\text{Strat}}$ $=\!\sum_{t=1}^{T}\xi_{t}^{\text{Strat}}\big(\sum_{1\leq t^{\prime}<t}G(t-t^{\prime})\xi_{t^{\prime}}^{\text{Strat}}\big)\!+\!\sum_{t=1}^{T}\frac{\xi_{t}^{\text{Strat}}s_{t}}{2}$ (35)
where we have used that $p_{t}=m_{t}+s_{t}/2$, with $s_{t}/2$ the half-spread. We have to note that in our modeling framework the sign of the metaorder is uncorrelated with the order flow from the rest of the market. We refer the reader to the following work [CHTH23] for an optimal trade execution framework with endogenous order flow, where a trader’s own order submissions trigger child orders and further influence future price dynamics. Equation (35) shows that the expected execution shortfall is partly due to the spread and partly due to the impacts of past trades. Our goal is to use the expected shortfall as our cost function to back-test the optimal execution strategy to realized market data, by minimizing this cost functional. We will also consider the simplified case where the transient market impact component boils down to the instantaneous impact only which we have estimated using the procedure described in Section 3.2. For the half-spread cost we use the empirical time-weighted half-spread directly obtained from the LOB data as described in Section 3.2.
We will implement the optimal execution strategy for each hourly product using the estimated parameters we have derived, both for the Hawkes process which models the arrival rates of trades (buy/sell). We are going to distinguish between the univariate and bivariate calibration using the obtained parameters we presented in Table 1 for the univariate case and Table 5 for the bivariate case accordingly. The transaction cost analysis results, will be devoted to each hourly product separately. Our goal is to test whether the optimal strategy allows for cost reduction when we are considering the liquidation of a fixed position of total $X_{0}=200-300$ MWh. The relative performance measures with respect to the optimal strategy is given by $r_{\text{Bench}}=(\mathcal{D}^{\text{Bench}}-\mathcal{D}^{\text{Strat}})/\mathcal{D}^{\text{Bench}}$. Starting from the univariate case, where the arrival of buy and sell market orders are independent, Figure 9 showcases the relative cost improvements (in percentage points) for each hourly trading product in comparison to the TWAP and VWAP strategies. The relative improvements in terms of cumulative execution costs are on average positive, indicating a clear cost reduction when the trader chooses to liquidate their position using the optimal execution strategy.
@@
Examining the average relative cost improvements associated with the optimal strategy and the two benchmarks for each maturity, we observe from Figure 11 and Table 3 that the optimal strategy achieves cost reduction for each product. Specifically, the cost reduction is very significant for the first hourly trading products (1h-4h) and then stabilizes for the rest of the trading hours, exhibiting a slight decrease for the mid-day trading products (12h-16h).
Considering the seasonality of liquidity across each trading product, as shown in Figure 2(a), there is a clear pattern in the associated liquidity of each hourly product. The most liquid products are associated with mid-day (noon) hours, while the first hourly trading products are the least liquid ones. This suggests that the amount of cost savings is negatively correlated with the average traded volume observed for each product. Therefore, comparatively less volatile, more liquid trading products at lower tick sizes may benefit less from the optimal trading strategy compared to the less liquid products, where the improvement is more pronounced.
Considering the time-dependent (pathwise, not only terminal) behavior of trading costs, Figure 10 displays the surface of the relative cost difference across the trading hours (last 8 trading hours in our backtests) and hourly trading product ($1,\dots,24$). By construction, positive values indicate that the benchmark incurs a higher cumulative cost than the optimal strategy at that point in time. For completeness, terminal (end-of-horizon) relative improvements are reported in Figures 11. In both benchmarks, the optimal strategy tends to incur smaller costs early in the session—consistent with its slower initial execution speed—and then accelerates as time to maturity shrinks. Immediately after the market decoupling into the local-trading regime, we observe a temporary dip in the mean relative improvement, reflecting wider bid–ask spreads and thinner liquidity; if an insufficient fraction of the position is completed before decoupling, the instantaneous cost increases. Nevertheless, the strategy recovers and the terminal relative improvements at maturity ($T=8$ h) remain positive against both TWAP and VWAP.
@@
Figure 11: Comparison of relative cost improvements across maturities.
Looking at the mean relative cost improvements of the bivariate model compared to the VWAP, TWAP and the univariate model Table 4 and Figure 12 showcase that the results compared to the benchmark strategies are again significant. However, compared to the univariate case, where we have only the self-excitation parameters in the exponential kernel, the results vary across each hourly trading product, indicating that there is no clear improvement when considering the full kernel for the modeling of the order flow of buy and sell MOs.
@@
Figure 12: Comparison of relative cost improvements across maturities under self- and cross-excitation.
@@ 5 Conclusion
There are several potential extensions to the methodology and modeling framework presented in this work. One possible extension is to model multiple contracts simultaneously rather than focusing on a single contract. Our current one dimensional optimal execution model does not capture the dependence between different maturities, which is crucial for valuing assets with payoffs depending on several maturities (e.g. battery valuation). This could be addressed by employing a multivariate transient price impact kernel which models how trades for a given trading hourly product would (directly or indirectly) impact the price of another product.
One could also formulate an optimal execution problem under a cross-impact propagator model, where in intraday markets overlapping sessions exist for different delivery days. Given the strong interdependencies among products, incorporating cross-product and cross-impact effects may offer deeper insights into price formation in intraday electricity markets.
@@ Acknowledgement
We sincerely thank the two anonymous referees for their constructive feedback, which greatly improved the clarity and presentation of this paper. Konstantinos Chatziandreou gratefully acknowledges the support through Statkraft Trading GmbH. Konstantinos Chatziandreou is grateful to Daniel Gruhlke and Simon Hirsch for many helpful discussions. This work contains the author’s opinion and does not necessarily reflect Statkraft’s position. The authors declare no conflict of interest.
@@ Appendix A MLE for Hawkes Parameters
$\lambda_{t}^{(p)}=\lambda_{\infty}^{(p)}+\sum_{m=1}^{P}\int_{-\infty}^{t}\phi_{p,m}(t-u)\,dN_{u}^{(m)},\quad p=1,\dots,P,$ (36)
@@
$\log\mathcal{L}\bigl(\Theta;\,\bm{\mathcal{T}}\bigr)=\sum_{p=1}^{P}\Big[\sum_{j=1}^{N_{T}^{(p)}}\log\lambda_{\,t_{j}^{(p)}}^{(p)}\int_{0}^{T}\lambda_{u}^{(p)}\,du\Big],$ (37)
@@ Appendix B Goodness of fit analysis
In this section we aim to describe how we selected the best model from the estimated point process models describing the order-flow. We will first introduce a general theory for goodness-of-fit analysis of point processes. Let the compensator (or the cumulative intensity process) of a point process $N_{t}$ be defined by the following integral of the conditional intensity
$\Lambda(t)=\int_{0}^{t}\lambda(v)dv$ (39)
By using the compensator, the technique of random time change transforms a point process with intensity $\lambda(t)$ into a standard Poisson process. The underlying rationale is that if the proposed point process model aligns well with the observed data, the time-changed process should exhibit the defining characteristics of a standard Poisson process.
Consider a point process $N_{t}$ whose events occur at times $\tau_{n}$. According to the theorem, the transformed arrival times $\tilde{\tau}_{n}=\Lambda(\tau_{n})$ correspond to the arrival times of a unit-rate Poisson process. As a result of the Papangelou’s theorem [daley2006introduction, Theorem 7.4.I] , the transformed interarrival times
$\tilde{\Delta}_{i}=\Lambda(\tau_{i+1})-\Lambda(\tau_{i})=\int_{\tau_{i}}^{\tau_{i+1}}\lambda(v)\,dv$ (40)
@@
B.1 Goodness of Fit Results: Comparison Piecewise and Spline approximation
@@
Starting from the univariate case, which is equivalent to the model where we have a zero component in the non-diagonals of the exponential kernel matrix of the Hawkes process we examine the two different choices of the baseline function. This model essentially boils down to the model where the order-flow is uncorrelated, meaning that buy market orders only excite orders of the same sign. The first and the most naive one would be to look into hourly blocks of the LOB and for each of these blocks to fit the underlying Hawkes process in order to account for the regime switch of the baseline intensity. Of course, this naive approach is also time-consuming. Considering a total of one year of back-test data, along with the calibration of each of the 24 products and the last 8 tradable hours, this approach does not seem appealing at first sight. Hence the second approach would be to consider an exponential baseline function. However, as we described above we are now going to consider a new approach by non-parametrically estimating the baseline intensity using the idea originally introduced in the paper [Chen_Hall_2013] to account for the empirical finding motivated by Figure 2(a), which describes the drop of trading activity during the local trading hours and the sharp increase towards delivery for the first trading hours. We can see from the following Figure 13 that both the piecewise approximation  [^1] Looking into hourly blocks of the LOB and fitting the Hawkes parameters by assuming a constant baseline intensity for each hourly block as well as the non-parametric approximation of the baseline intensity seem to capture the increase of the market activity for both the sell and buy side of the LOB. However, it is clear from the graph that the median baseline intensities estimated using the non-parametric approach seem to better fit the drop in the market activity immediately after the start of the local trading regime, for both sides of the LOB; which is also in agreement with the empirical intensities presented in Figure 2(a). In Figure 13 we showcase the result only for the hourly products (16h - 18h). The total number of basis points used for the estimation of the baseline intensity is $N_{basis}=10$. One more nice feature that we observe is that on average the shape of the baseline function for each delivery hour seems to be close. This is also showcased in Figure 4 where we present the calibrated baseline intensity for each side of the LOB and each hourly product.
@@
In order to evaluate the performance of each model we perform a Goodness-of Fit analysis. For the comparison of the models, we use the calibrated models for each maturity (24 in total) between the dates (2023-01-26 until 2023-12-03 a total of 311 days). Hence we have a total of 7464 models for each different approximation of the baseline intensity. Since both models have been calibrated using a Maximum Likelihood method, one way to compare the fitting performance of each model is by using a Likelihood ratio test. However, in this work, we simply use a more informal inspection of the differences in the maximum log-likelihood values of the model with a baseline piecewise constant function and the model with a spline approximation of the baseline function. Since we are calibrating each model independently we showcase the differences for each hourly contract and each daily calibration. The following Figure showcases the box plots of the difference for each hourly product and each orderbook side (Buy/Sell). For both sides and each hourly the Maximized value of the log-likelihood function for the model with piece wise constant baseline function is on average lower than the spline approximation model, which is attributed to the positive differences of the maximum likelihood values showcased in the Box plots 15 - 16  [^2] We note that the boxplots include all fitted models; (2023-01-26 until 2023-12-03 a total of 311 days) for buy and sell market orders respectively..
@@
In a similar fashion we compare the AIC differences for each hourly product. Recall that in comparison to the log-likelihood tests the AIC scores indicate the best-performing model when the score is smaller. From the following figure, we can again see that the AIC score of the spline model is smaller (since we are testing the differences between (AIC Piecewise- AIC Spline) and we can see a positive difference between the two) for each calibrated trading hour. This indicates that the performance of the spline fit on the baseline intensity of the Hawkes process provides a better fit to the data. This result is obtained for both sides of the LOB (Buy/Sell) (See Figures 17-18).
@@
Since we have multiple calibrated models for each day and each maturity using a QQ plot between the compensated interarrival times and a unit exponential distribution would not be a very informative way to derive a qualitative conclusion on the performance of each separate model. Hence we choose to use performance metrics instead as the Anderson and Darling test for the exponential distribution and the Kolmogorov-Smirnov (KS) test. We note that in general the Anderson-Darling test is much more sensitive to the tails of the distribution, whereas KS test is more aware of the center of the distribution. One more way to test the "distance" of the time-changed inter-arrival times and the unit exponential distribution is to use a distance function defined between probability distributions on a given metric space. Suppose $\mathbb{P}$ and $\mathbb{Q}$ are two probability distributions on the real line, with corresponding cumulative distribution functions $F$ and $G$. The Wasserstein distance $W(\mathbb{P},\mathbb{Q})$ is defined as $W(\mathbb{P},\mathbb{Q})=\int_{-\infty}^{\infty}|F(x)-G(x)|\,dx.$
This enables us to validate the estimated model using both the Anderson-Darling Test as well as the Wasserstein distance on an exponential distribution with parameter $\lambda=1$ (see also [Bowsher2007]). The following figures depict the Wasserstein distances for buy and sell market orders respectively and include once more all the fitted models (2023-01-26 until 2023-12-03). We only showcase the results for three maturities due to a lack of space, however, the rest of the calibrated hourly products can be found in Appendix C. Looking at the Buy/Sell side, the median Wasserstein distance for the model with the spline approximation of the baseline intensity is significantly lower than for the piece wise approximation (see Figures 19-20). Additionally, looking at the inter quantile range (IQR) of the distance for the case of the spline model, it is considerably lower than the piecewise approximation. The Anderson-Darling test showcases that the fit to a unit exponential is better on both sides of the LOB since the values of the test statistic are lower in both cases. Finally, we examine the independence of the time-changed interarrival times using the Ljung-Box test. The results of the p-values plot in Figures 23-24, again reveal that since the mean of the p-values of all fitted models is not below $0.05$, which is the selected significance level, both tests for the spline and piecewise models agree that you can not reject the null of no auto-correlation between the series and each of its first 100 lags with > 95% confidence level. However, testing the number of times that the null hypothesis was rejected between all fitted models (days), and examining the difference between the spline and the piece wise model, we can conclude that this number is smaller for the spline model, indicating that most of the time the time-changes inter arrival times of the spline model are indeed serially uncorrelated.